# Conformalized Polynomial Chaos Expansion for Uncertainty-aware Surrogate Modeling

### v1

@@ Abstract
This work introduces a method to equip data-driven polynomial chaos expansion surrogate models with intervals that quantify the predictive uncertainty of the surrogate. To that end, we integrate jackknife-based conformal prediction into regression-based polynomial chaos expansions. The jackknife algorithm uses leave-one-out residuals to generate predictive intervals around the predictions of the polynomial chaos surrogate. The jackknife+ extension additionally requires leave-one-out model predictions. The key to efficient implementation is to leverage the linearity of the polynomial chaos regression model, so that leave-one-out residuals and, if necessary, leave-one-out model predictions can be computed with analytical, closed-form expressions, thus eliminating the need for repeated model re-training. In addition to the efficient computation of the predictive intervals, a significant advantage of this approach is its data efficiency, as it requires no hold-out dataset for prediction interval calibration, thus allowing the entire dataset to be used for model training. The conformalized polynomial chaos expansion method is validated on several benchmark models, where the impact of training data volume on the predictive intervals is additionally investigated.
@@ 1.  Introduction
Originally developed as an efficient uncertainty quantification (UQ) method [1, 2, 3], the polynomial chaos expansion (PCE) is in recent years routinely employed as a machine learning (ML) regression model [4]. Within this context, regression-based PCEs have been used as data-driven surrogate models in various science and engineering applications [5, 6, 7, 8, 9]. Their use has also been extended to scientific ML tasks like manifold learning [10, 11, 12, 13], operator learning [14], and physics-informed ML [15, 16].
@@
In contrast to the use of PCEs for UQ purposes, PCE-based data-driven predictive models are purely deterministic. That is, they yield point predictions that lack any type of uncertainty or confidence metrics, as opposed to Gaussian processes (GPs) [17] or Bayesian neural networks (NNs) [18], for example. Ensemble modeling and bootstrapping are common ways of providing such predictive uncertainty intervals to data-driven models, including PCEs [19, 20]. Other relevant approaches include combined PCE-GP models [21, 22] or stochastic PCE formulations [23]. However, these approaches induce either specific modeling limitations or significant computational overhead due to model retraining.
@@
Considering predictive UQ for ML models, conformal prediction (CP) [24] is an established method for producing statistically valid prediction intervals given a point predictor, only under the assumption of data exchangeability. The resulting predictive models are commonly called “conformalized” [25, 26, 27, 28, 29]. CP is a distribution-free UQ method, meaning that it does not require any assumptions about the underlying data distribution or the structure of the data-driven model. Additionally, certain CP methods provide statistically rigorous, finite-sample guarantees regarding prediction interval coverage. On its downsides, CP does not yield full probability distributions like Bayesian methods do. Moreover, its coverage is marginal instead of conditional, hence possibly under- or overconfindent in different regions of the input space. Adaptive CP methods that address the latter issue are actively pursued [30, 31]. Data demand can also be an issue, in particular considering so-called inductive or split CP methods [32, 33], which require that part of the available data is excluded from training and reserved for prediction interval estimation. Remedies have been sought in so-called cross CP approaches leveraging cross-validation techniques [34, 35], including jackknife-based variants that utilize leave-one-out (LOO) residuals [33, 36]. Similar to bootstrapping and ensemble modeling, the downside of these methods is the need for repeated model retraining, which can result in undesired computational overhead.
@@
This work proposes a novel combination of regression-based PCEs with CP based on the jackknife method [33] and its jackknife+ extension [36], for constructing predictive intervals that quantify the predictive uncertainty of data-driven PCEs. Combining CP with data-driven surrogate models has been explored before, e.g., for GPs [29] or various NNs architectures [37]. However, the use of regression-based PCEs results in an uncertainty-aware, data-driven surrogate modeling framework, which is additionally very computationally efficient. The key enabler is the linear form of PCE regression models, which allows to compute the LOO residuals required by the jackknife method analytically in closed form, without the need to re-train the PCE. The same is true for the LOO model predictions needed by the jackknife+ method. Due to this synergy of jackknife-based CP and PCE regression, the conformalized PCE method provides predictive intervals with minimal computational overhead. Additionally, the suggested approach belongs to the class of cross CP methods, thus allowing the full available dataset to be used for training. This is a crucial advantage in small-data learning regimes. To the authors’ knowledge, this is the first work to develop a conformalized PCE regression method.
@@
The PCE method is implemented for total-degree PCEs utilizing ordinary least squares (OLS) for the computation of the PCE coefficients, and validated on four different benchmark models with one to ten inputs and scalar outputs. Extensions to other types of polynomial bases and coefficient estimation methods are conceptually straightforward, with the requirement for a fixed polynomial basis so that the closed-form formulas for the LOO residuals and model predictions remain valid (see section 2.3). Different training data availability conditions are also examined regarding their impact on prediction intervals and coverage. Prior works have shown the importance of training data volume for the accuracy and stability of regression-based PCEs [38, 39]. Here, we extend this investigation to assess how data availability impacts the quality of the predictive UQ provided by the conformalized PCE method.
@@
The remaining of this paper is organized as follows. Section 2 presents the methodology behind the suggested method, first recalling the necessary prerequisites of PCEs and CP, and then presenting the conformalized PCE. The method is validated using several benchmark models in section 3. The paper concludes with section 4, which summarizes the findings of this work and discusses follow-up research opportunities.
@@ 2.  Methodology
We begin by presenting the PCE regression model in section 2.1, followed by a presentation of CP section 2.2. The latter emphasized the jackknife and jackknife+ CP methods. Section 2.3 shows how LOO residuals and model predictions of PCE regression models can be computed with analytical, closed-form expressions, as to enable the cost-effective estimation of predictive intervlas based on jackknife and jackknife+ CP. Last, section 2.4 discusses non-conformity score normalization for conformalized PCEs.
@@ 2.1 Regression-based polynomial chaos expansion
In the following we consider a computational model, denoted as $\mu(\mathbf{x})$, which receives an input vector $\mathbf{x}\in\mathbb{R}^{N}$ and computes a scalar output $y\in\mathbb{R}$, such that $y=\mu\left(\mathbf{x}\right)$. The scalar output is considered for simplicity only; vector-valued outputs can also be considered by an element-wise application of all operations described next. The form of the model can range from a simple analytical function to a complex numerical solver. The only assumption is that the model is deterministic, i.e., the same output $y$ is observed given the same input $\mathbf{x}$.
@@
Next, we assume that the random realizations $\mathbf{x}=\mathbf{X}\left(\omega\right)$ coincide with the inputs of the computational model $\mu(\mathbf{x})$ and that the Doob-Dynkin lemma is satisfied [40]. Then, by propagation of uncertainty, the model’s response is a random variable dependent on the input random vector, i.e., $Y=\mu\left(\mathbf{X}\right)$. Provided that its variance is finite, the random output can be represented in the form of a PCE, such that
@@
where $\Lambda\subset\mathbb{N}_{0}^{N}$ is a multi-index set with cardinality $\Lambda=K$. Each index $k=1,\dots,K$ corresponds to a unique multi-index $\boldsymbol{\alpha}\in\Lambda$. One common truncation method is to use a so-called total-degree basis, where $\Lambda=\left\{\boldsymbol{\alpha}:\left|\boldsymbol{\alpha}\right|_{1}\leq P\right\}$ and $P$ denotes the maximum polynomial degree. The size of the total-degree basis is equal to $K=\frac{\left(N+P\right)!}{N!P!}$, which can become intractable for high polynomial degrees or high-dimensional inputs, the latter being a manifestation of the so-called curse of dimensionality. To alleviate this problem, several methods have been suggested for constructing sparse bases [41], e.g., by means of dimension-adaptive algorithms [42].
Given a multi-index set $\Lambda$ and the corresponding polynomial basis, the remaining task is to estimate the PCE coefficients $\mathbf{c}=\left(c_{\boldsymbol{\alpha}}\right)_{\boldsymbol{\alpha}\in\Lambda}^{\top}\in\mathbb{R}^{K\times 1}$. We opt here for a regression-based approach [38, 39, 43]. Alternatives include include pseudo-spectral projection [44, 45] or interpolation [46, 47, 48]. For the regression-based coefficient estimation, we assume a set of input parameter realizations along with the corresponding model responses $\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$, commonly called the experimental design or, in an ML-context, the training data set. We then construct the so-called design matrix $\mathbf{D}\in\mathbb{R}^{M\times K}$, such that $d_{mk}=\Psi_{k}\left(\mathbf{x}^{(m)}\right)$. Collecting the model responses in a vector $\mathbf{y}=\left(y^{(1)},\dots,y^{(M)}\right)^{\top}\in\mathbb{R}^{M\times 1}$, the coefficients can be computed by solving the least squares minimization problem
@@
For the regression problem (4) to be well posed, the experimental design must be at minimum equal to the polynomial basis. Larger experimental designs are typically needed for well determined problems. In fact, the size of the experimental design is crucial for the stability and accuracy of the resulting PCE [38, 39]. This limitation can be overcome with sparse PCE methods, where the problem (4) is complemented with a penalty term that induces sparsity in its solution. Common techniques include the least absolute shrinkage and selection operator (LASSO) and compressive sensing [41].
@@ 2.2 Conformal prediction
such that $y^{*}\in\widehat{C}_{1-s}\left(\mathbf{x}^{*}\right)$ implies that $\alpha\!\left(\mathbf{x}^{*},y^{*};\widehat{\mu}\right)\leq\widehat{q}_{1-s}\!\left\{\alpha_{j}\right\}$ and vice versa. One common non-conformity score is the absolute residual, such that $\alpha_{j}=\left|r_{j}\right|=\left|y^{(j)}-\widehat{\mu}\left(\mathbf{x}^{(j)}\right)\right|$. This choice results in the construction of the symmetric predictive interval
@@
A naive approach to estimate the predictive interval (8) would be to use the residuals on the training data $\mathcal{D}_{\text{train}}=\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$. However, training data residuals are expected to be smaller compared to residuals obtained from previously unseen test data points. Hence, the resulting predictive interval will most probably not provide the required coverage. Split or inductive CP approaches [32, 33] use a separate validation dataset $\mathcal{D}_{\text{val}}\neq\mathcal{D}_{\text{train}}$ to compute the non-conformity scores and construct the prediction interval (8). Split CP guarantees the required finite-sample coverage (6), however, this comes at the cost of larger data demand due to the need for a validation dataset. To avoid the need for additional data, cross CP [34, 35] utilizes $K$-fold cross-validation, such that the training dataset $\mathcal{D}_{\text{train}}$ is split into $K$ disjoint subsets $\mathcal{D}_{1},\dots,\mathcal{D}_{K}$ of (approximately) equal size. Then, the regression models $\widehat{\mu}_{\sim k}\left(\mathbf{x}\right)$ are trained using the datasets $\mathcal{D}_{\sim k}=\mathcal{D}_{\text{train}}\setminus\mathcal{D}_{k}$, $k=1,\dots,K$, and the cross-validation residuals
@@
A special case of cross-validation is the jackknife method [49]. Using the jackknife in the context of CP [33], the non-conformity scores are computed using the LOO residuals
@@
where the regression model $\widehat{\mu}_{\sim m}\left(\mathbf{x}\right)$ is trained using the dataset $\mathcal{D}_{\sim m}=\mathcal{D}_{\text{train}}\setminus\left\{\left(\mathbf{x}^{(m)},y^{(m)}\right)\right\}$. Using the absolute LOO residuals as non-conformity scores, the predictive interval (8) takes two equivalent forms, i.e.,
@@
Being a cross CP variant, jackknife CP also yields an empirical coverage close to the target level, also without formal coverage guarantees. Barber et al. [36] suggested an extension of jackknife CP, called jackknife+, where the predictive interval is constructed as
@@
In contrast to the jackknife interval (11) which is centered at the prediction $\widehat{\mu}\left(\mathbf{x}^{*}\right)$ of the fully trained model, the jackknife+ interval (12) is based on the ensemble of the LOO predictions $\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)$, $m=1,\dots,M$. Hence, the jackknife+ predictive interval is asymmetric. This slight modification results in a formal coverage guarantee of $\left(1-2s\right)$, while the empirical coverage remains close to the target level (1s)(1-s), similar to cross CP and jackknife CP.
It should be noted that, in practice, the jackknife and jackknife+ CP methods produce almost identical predictive intervals and both achieve an empirical coverage close to the target level (1s)(1-s) [36]. Moreover, both jackknife-based approaches entail training $M$ regression models, thus being the extreme case of cross CP in terms of computational cost. However, for linear regression models like the PCE, the LOO residuals $r_{m}^{\text{LOO}}$ and the LOO predictions $\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)$ can be obtained without model retraining, as shown in section 2.3.
2.3 Conformalized polynomial chaos expansion
We now assume that $\widehat{\mu}\left(\mathbf{x}\right)$ is a regression-based PCE in the form of (3) and trained upon a dataset $\mathcal{D}_{\text{train}}=\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$. Then, the predictive intervals (11) (for jackknife) and (12) (for jackknife+) can be estimated without the need to retrain the PCE $M$ times.
Starting with jackknife method, only the LOO residuals defined in (10) are needed to estimate the prediction interval (11). Recalling from section 2.1 the design matrix $\mathbf{D}\in\mathbb{R}^{M\times K}$, we define the so-called hat matrix $\mathbf{H}\in\mathbb{R}^{M\times M}$ as

### v2

@@ Abstract
This work introduces a method to equip data-driven polynomial chaos expansion surrogate models with intervals that quantify the predictive uncertainty of the surrogate. To that end, jackknife-based conformal prediction is integrated into regression-based polynomial chaos expansions. The jackknife algorithm uses leave-one-out residuals to generate predictive intervals around the predictions of the polynomial chaos surrogate. The jackknife+ extension additionally requires leave-one-out model predictions. Both methods allow to use the entire dataset for model training and do not require a hold-out dataset for prediction interval calibration. The key to efficient implementation is to leverage the linearity of the polynomial chaos regression model, so that leave-one-out residuals and, if necessary, leave-one-out model predictions can be computed with analytical, closed-form expressions. This eliminates the need for repeated model re-training. The conformalized polynomial chaos expansion method is first validated on four benchmark models and then applied to two electrical engineering design use-cases. The method produces predictive intervals that provide the target coverages, even for low-accuracy models trained with small datasets. At the same time, training data availability plays a crucial role in improving the empirical coverage and narrowing the predictive interval, as well as in reducing their variability over different training datasets.
@@ 1.  Introduction
Originally developed as an efficient uncertainty quantification (UQ) method [spanos1989stochastic, xiu2002wiener, ghanem2003stochastic], the polynomial chaos expansion (PCE) is in recent years routinely employed as a machine learning (ML) regression model [torre2019data]. Within this context, regression-based PCEs are often used as data-driven surrogate models in various science and engineering applications [haghi2022surrogate, loukrezis2022power, campos2023polynomial, manfredi2024nonparametric, dai2025cost, lee2025surrogate]. Their use has also been extended to scientific ML tasks like manifold learning [soize2017polynomial, kontolati2022manifold, kontolati2022survey, giovanis2024polynomial], operator learning [sharma2025polynomial], and physics-informed ML [novak2024physics, sharma2024physics]. In contrast to the use of PCEs for UQ purposes, PCE-based data-driven predictive models are purely deterministic. That is, they yield point predictions that lack any type of uncertainty or confidence metrics. Models that provide such metrics are, for example, Gaussian processes (GPs) [williams2006gaussian] and Bayesian neural networks (NNs) [jospin2022hands]. Ensemble modeling and bootstrapping are common ways of providing such predictive uncertainty intervals to data-driven models, including PCEs [marelli2018active, shang2024active]. Other relevant approaches include combined PCE-GP models [schobi2015polynomial, kersaudy2015new] or stochastic PCE formulations [zhu2023stochastic]. However, these approaches induce either specific modeling limitations or significant computational overhead due to model retraining.
@@
Considering predictive UQ for ML models, conformal prediction (CP) [fontana2023conformal] is an established method for producing statistically valid prediction intervals given a point predictor, only under the assumption of data exchangeability. The resulting predictive models are commonly called “conformalized” [burnaev2016conformalized, romano2019conformalized, huang2023uncertainty, podina2024conformalized, jaber2025conformal]. CP is a distribution-free UQ method, meaning that it does not require any assumptions about the underlying data distribution or the structure of the data-driven model. Additionally, certain CP methods provide statistically rigorous, finite-sample guarantees regarding prediction interval coverage. On its downsides, CP does not yield full probability distributions like Bayesian methods do. Moreover, its coverage is marginal instead of conditional, hence possibly under- or overconfindent in different regions of the input space. Adaptive CP methods that address the latter issue are actively pursued [zaffran2022adaptive, deutschmann2024adaptive]. Data demand can also be an issue, in particular considering so-called inductive or split CP methods [papadopoulos2002inductive, lei2018distribution], which require that part of the available data is excluded from training and reserved for prediction interval estimation. Remedies have been sought in so-called cross CP approaches leveraging cross-validation techniques [vovk2015cross, vovk2018cross], including jackknife-based variants that utilize leave-one-out (LOO) residuals [lei2018distribution, barber2021predictive]. Similar to bootstrapping and ensemble modeling, the downside of these methods is the need for repeated model retraining, which can result in undesired computational overhead.
@@
This work proposes a novel combination of regression-based PCEs with CP based on the jackknife method [lei2018distribution] and its jackknife+ extension [barber2021predictive]. The resulting conformalized PCE models are equipped with predictive intervals that quantify the uncertainty in their predictions. Combining CP with data-driven surrogate models has been explored before, e.g., for GPs [jaber2025conformal] or various NNs architectures [gopakumar2024uncertainty]. However, the use of regression-based PCEs results in an uncertainty-aware, data-driven surrogate modeling framework, which is additionally very computationally efficient. The key enabler is the linear form of PCE regression models, which allows to compute the LOO residuals required by the jackknife method analytically in closed form, without the need to re-train the PCE. The same is true for the LOO model predictions needed by the jackknife+ method. Due to this synergy of jackknife-based CP and PCE regression, the predictive intervals are estimated with minimal computational overhead. The suggested approach belongs to the class of cross CP methods, thus allowing the full available dataset to be used for training. This is a crucial advantage in small-data learning regimes. To the authors’ knowledge, this is the first work to develop a conformalized PCE regression method.
@@
The method is implemented for total-degree PCEs utilizing ordinary least squares (OLS) for the computation of the PCE coefficients. Extensions to other types of polynomial bases and coefficient estimation methods are conceptually straightforward, with the requirement for a fixed polynomial basis so that the closed-form expressions for the LOO residuals and model predictions remain valid (see section 2.3). The method is first validated on four benchmark models and then applied to two electrical engineering design use-cases. The first use-case concerns the design of a heat sink for power electronics cooling [loukrezis2022power]. The second use-case concerns the design of a Stern-Gerlach electromagnet [masschaele2011design]. In all numerical experiments, we assess whether the conformalized PCE surrogate provides predictive uncertainty intervals that result in empirical coverages similar to the required target levels. Additionally, we investigate the impact of different conformalized PCE configurations, with an emphasis on training data availability.
@@
The remaining of this paper is organized as follows. Section 2 presents the methodology behind the suggested method, first recalling the necessary prerequisites of PCE and CP methods, and then presenting the conformalized PCE. The numerical experiments are presented in section 3. Section 4 summarizes the findings of this work and discusses follow-up research opportunities.
@@ 2.  Methodology
We begin by presenting the PCE regression model in section 2.1, followed by a presentation of CP section 2.2. The latter focuses on the jackknife and jackknife+ CP methods. Section 2.3 shows how LOO residuals and predictions of PCE regression models can be computed with analytical, closed-form expressions, as to enable the cost-effective estimation of predictive intervlas based on jackknife and jackknife+ CP. Last, section 2.4 discusses non-conformity score normalization for conformalized PCEs.
@@ 2.1 Regression-based polynomial chaos expansion
In the following we consider a computational model, denoted as $\mu(\mathbf{x})$, which receives an input vector $\mathbf{x}\in\mathbb{R}^{N}$ and computes a scalar output $y\in\mathbb{R}$, such that $y=\mu\left(\mathbf{x}\right)$. While scalar outputs are considered throughout this work, vector-valued outputs can also be considered by an element-wise application of all operations described next. The form of the model can range from a simple analytical function to a complex numerical solver. The only assumption is that the model is deterministic, i.e., the same output $y$ is observed given the same input $\mathbf{x}$.
@@
Next, we assume that the random realizations $\mathbf{x}=\mathbf{X}\left(\omega\right)$ coincide with the inputs of the computational model $\mu(\mathbf{x})$ and that the Doob-Dynkin lemma is satisfied [rao2006probability]. Then, by propagation of uncertainty, the model’s response is a random variable dependent on the input random vector, i.e., $Y=\mu\left(\mathbf{X}\right)$. Provided that its variance is finite, the random output can be represented in the form of a PCE, such that
@@
where $\Lambda\subset\mathbb{N}_{0}^{N}$ is a multi-index set with cardinality $\Lambda=K$. Each index $k=1,\dots,K$ corresponds to a unique multi-index $\boldsymbol{\alpha}\in\Lambda$. One common truncation method is to use a so-called total-degree basis, where $\Lambda=\left\{\boldsymbol{\alpha}:\left|\boldsymbol{\alpha}\right|_{1}\leq P\right\}$ and $P$ denotes the maximum polynomial degree. The size of the total-degree basis is equal to $K=\frac{\left(N+P\right)!}{N!P!}$ and becomes intractable for high polynomial degrees or input dimensions. To alleviate this problem, several methods have been suggested for constructing sparse bases [luethen2021sparse], e.g., by means of dimension-adaptive algorithms [luethen2022automatic].
Given a multi-index set $\Lambda$ and the corresponding polynomial basis, the remaining task is to estimate the PCE coefficients $\mathbf{c}=\left(c_{\boldsymbol{\alpha}}\right)_{\boldsymbol{\alpha}\in\Lambda}^{\top}\in\mathbb{R}^{K\times 1}$. We opt here for a regression-based approach [migliorati2013approximation, migliorati2014analysis, hadigol2018least]. Alternatives include include pseudo-spectral projection [constantine2012sparse, conrad2013adaptive] or interpolation [buzzard2012global, porta2014inverse, galetzka2023hp]. For the regression-based coefficient estimation, we assume a set of input parameter realizations along with the corresponding model responses $\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$, commonly called the experimental design or, in an ML-context, the training data set. We then construct the so-called design matrix $\mathbf{D}\in\mathbb{R}^{M\times K}$, such that $d_{mk}=\Psi_{k}\left(\mathbf{x}^{(m)}\right)$. Collecting the model responses in a vector $\mathbf{y}=\left(y^{(1)},\dots,y^{(M)}\right)^{\top}\in\mathbb{R}^{M\times 1}$, the coefficients can be computed by solving the least squares minimization problem
@@
For the regression problem (4) to be well posed, the experimental design must be at minimum equal to the polynomial basis. Larger experimental designs are typically needed for well determined problems. In fact, the size of the experimental design is crucial for the stability and accuracy of the resulting PCE [migliorati2013approximation, migliorati2014analysis]. This limitation can be overcome with sparse PCE methods, where the problem (4) is complemented with a penalty term that induces sparsity in its solution. Common techniques include the least absolute shrinkage and selection operator (LASSO) and compressive sensing [luethen2021sparse].
@@ 2.2 Conformal prediction
such that $y^{*}\in\widehat{C}_{1-s}\left(\mathbf{x}^{*}\right)$ implies that $\alpha\!\left(\mathbf{x}^{*},y^{*};\widehat{\mu}\right)\leq\widehat{q}_{1-s}\!\left\{\alpha_{j}\right\}$ and vice versa. A commonly used non-conformity score is the absolute residual, such that $\alpha_{j}=\left|r_{j}\right|=\left|y^{(j)}-\widehat{\mu}\left(\mathbf{x}^{(j)}\right)\right|$. This choice results in the construction of the symmetric predictive interval
@@
A naive approach to estimate the predictive interval (8) would be to use the residuals on the training data $\mathcal{D}_{\text{train}}=\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$. However, training data residuals are expected to be smaller compared to residuals obtained from previously unseen test data points. Hence, the resulting predictive interval will most probably not provide the required coverage. Split or inductive CP approaches [papadopoulos2002inductive, lei2018distribution] use a separate validation dataset $\mathcal{D}_{\text{val}}\neq\mathcal{D}_{\text{train}}$ to compute the non-conformity scores and construct the prediction interval (8). Split CP guarantees the required finite-sample coverage (6), however, this comes at the cost of a larger data demand. To avoid the need for additional data, cross CP [vovk2015cross, vovk2018cross] utilizes $K$-fold cross-validation, such that the training dataset $\mathcal{D}_{\text{train}}$ is split into $K$ disjoint subsets $\mathcal{D}_{1},\dots,\mathcal{D}_{K}$ of (approximately) equal size. Then, the regression models $\widehat{\mu}_{\sim k}\left(\mathbf{x}\right)$ are trained using the datasets $\mathcal{D}_{\sim k}=\mathcal{D}_{\text{train}}\setminus\mathcal{D}_{k}$, $k=1,\dots,K$, and the cross-validation residuals
@@
A special case of cross-validation is the jackknife method [chernick2012jackknife]. Using the jackknife in the context of CP [lei2018distribution], the non-conformity scores coincide with the absolute values of the LOO residuals
@@
where the regression model $\widehat{\mu}_{\sim m}\left(\mathbf{x}\right)$ is trained using the dataset $\mathcal{D}_{\sim m}=\mathcal{D}_{\text{train}}\setminus\left\{\left(\mathbf{x}^{(m)},y^{(m)}\right)\right\}$. Then, the predictive interval (8) takes two equivalent forms, i.e.,
@@
Being a cross CP variant, jackknife CP also yields an empirical coverage close to the target level, also without formal coverage guarantees. Barber et al. [barber2021predictive] suggested an extension of jackknife CP, called jackknife+, where the predictive interval is constructed as
@@
In contrast to the jackknife interval (11) which is centered at the prediction $\widehat{\mu}\left(\mathbf{x}^{*}\right)$ of the fully trained model, the jackknife+ interval (12) is based on the ensemble of the LOO predictions $\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)$, $m=1,\dots,M$. Hence, the jackknife+ interval (12) is asymmetric. This slight modification results in a formal coverage guarantee of $\left(1-2s\right)$, while the empirical coverage remains close to the target level (1s)(1-s), similar to cross CP and jackknife CP.
In practice, both the jackknife and jackknife+ CP methods achieve an empirical coverage close to the target level (1s)(1-s) and produce almost identical predictive intervals [barber2021predictive]. Moreover, both approaches entail training $M$ regression models, thus being extreme cases of cross CP in terms of computational cost. However, for linear regression models like the PCE, the LOO residuals $r_{m}^{\text{LOO}}$ and the LOO predictions $\widehat{\mu}_{\sim m}\left(\mathbf{x}^{*}\right)$ can be obtained without model retraining, as shown in section 2.3.
2.3 Conformalized polynomial chaos regression
We now assume that $\widehat{\mu}\left(\mathbf{x}\right)$ is a regression-based PCE in the form of (3) and trained with a dataset $\mathcal{D}_{\text{train}}=\left\{\mathbf{x}^{(m)},y^{(m)}\right\}_{m=1}^{M}$. Then, the predictive intervals (11) (for jackknife) and (12) (for jackknife+) can be estimated without the need to retrain the PCE $M$ times.
Starting with the jackknife method, only the LOO residuals defined in (10) are needed to estimate the prediction interval (11). Recalling from section 2.1 the design matrix $\mathbf{D}\in\mathbb{R}^{M\times K}$, the so-called hat matrix $\mathbf{H}\in\mathbb{R}^{M\times M}$ is defined as